# The Study of the Canonical forms of Killing tensor in vacuum with $\Lambda$

### v1

@@ 4.2 Canonical forms
Obtaining the canonical forms of a symmetric 2nd-rank tensor can be a challenging task when approached algebraically. However, in a symmetric matrix the additional symmetries could aid us to find its canonical forms geometrically. The canonical forms contain the minimum number of independent scalars which compose the eigenvalues during the diagonalization [Churchill, 1932]. The equation for diagonalization in any tensor takes the following form.
${K^{\mu}}_{\nu}z^{\nu}-\lambda z^{\mu}=0$ (29)
In this manner, it becomes clear that the operation of our tensor or our linear vector function on a vector leaves it unchanged. In this context, it is known that every 2nd-order symmetric tensor defines a linear mapping that transforms a vector k into another vector v unless the vector is an eigenvector. We are searching for these types of directions, which, of course, do not alter the quadratic form of the metric. In our formalism, the norm of a vector must remain invariant.
@@
Algebraically, the result is that a symmetric matrix $K$ is reducible under an orthogonal transformation with a matrix P to a canonical form $PKP^{-1}$ in which all non-diagonal elements are zero.
@@
A similar analysis can also be found in Landau’s book on p. 271 regarding the stress-energy-momentum tensor [Landau and Lifschitz, 1975]. In this section, he mentions that this procedure is, in fact, the application of the “Principal Axis Theorem” or “Spectral Theorem” for a matrix. However, it appears that the author applies diagonalization to a symmetric tensor in a covariant form. This approach is not correct in general. Diagonalization should be applied to a tensor in a mixed tensor form, as shown in equation (29). This is always valid because a term proportional to the metric merely shifts all eigenvalues by the same amount, the latter is also denoted in [Stephani et al., 2009].
4.2.1 The presence of null vectors within planes
In line with the work of Churchill we obtained the Canonical forms of Killing tensor [Churchill, 1932]. The study of Churchill was operated in pseudo-Euclidean spacetime with signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and the object of his study was a linear symmetric vector function whose components are represented by a symmetric tensor of valence 2. In this reference Churchill was based on the work of Rainich who studied the antisymmetricity of the electromagnetic tensor [Rainich, 1925] and he remarks that: “It is known that every linear vector function in four-dimensional space has at least one invariable plane”.
Due to the last statement Churchill operated his calculation with a specific manner. He lied his $k,l$ vectors (the corresponding $m,\bar{m}$ of our formalism) in the invariable plane which does not contain any null vectors. With this choice all the canonical forms has the same downward-right block where there are two real distinct eigenvalues. Hence, all forms have different eigenvalues only in the upward-left block where the existence of null vectors takes place.
In the following segment we present the classification of canonical forms, therein the statement of the previous paragraph is evident since all the canonical forms proved to have the same eigenvalues $-(\lambda_{2}\pm\lambda_{7})$ in the 2-dimensional plane where non-null vectors are present. Although, the $K^{0}$ form serves as an exception since it has one triple eigenvalue with multiplicity 1. Due to this necessary condition one can prove that the triple eigenvalue satisfies either the first relation
$\lambda_{1}=-(\lambda_{2}+\lambda_{7})$
or the second one.
$\lambda_{1}=-(\lambda_{2}-\lambda_{7})$
After this introduction let us proceed to the following classification. Our pseudo-Euclidean spacetime comprises three types of planes [Hall, 1976]. It was helpful to mention that our framework is described by a null tetrad frame, hence, our forms have a totally different shape. Besides, we categorize our cases differently. In our Case 0, there is a plane with one null vector and it is characterized as a singular case. In Case 1 and Case 3, there are two null vectors lie within the plane, and in Case 2, there are no null vectors within our plane [^5] For this case we expect that the two blocks upward-left and downward-right have the exact same form with different components. [Rainich, 1952].
In Case 0 there is only one null vector lies within the plane, using this we obtain the $K^{0}$ form. Using the diagonalization procedure one can easily find that it has one triple eigenvalue and it is a Jordan canonical form. The diagonalization reveals that in order to have one timelike and three spacelike eigenvectors as a necessary condition for our canonical form there are two cases for the triple eigenvalue as we mentioned few lines above. Thus, the canonical form for $K^{0}$ results to
${K{{}^{0}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&-p&-\bar{p}\\ \lambda_{1}&0&0&0\\ -p&0&\lambda_{7}&\lambda_{2}\\ -\bar{p}&0&\lambda_{2}&\lambda_{7}\end{pmatrix}\hskip 8.5359ptp=-\bar{p}=\pm 1$ (31)
@@ 4.2.1 The presence of null vectors within planes
In Case 1 we have two canonical forms with one double eigenvalue called as $K^{1}$. This case produces two forms but considered as one by us. The reason is based on the symmetry of the symmetrical null tetrad frame. This symmetry referred to the interchanges between the tetrads $n^{\mu}\leftrightarrow l^{\mu}$ and $m^{\mu}\leftrightarrow\bar{m}^{\mu}$. Regarding this, even if the spin coefficients interchange the result remains the same.
${K{{}^{1a}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}\hskip 2.84544pt{K{{}^{1b}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&0&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (32)
Case 2 contains also a canonical form with 4 distinct real eigenvalues called as $K^{2}$.
${K{{}^{2}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (33)
Finally, when considering the $K^{3}$ form, it possesses a pair of eigenvalues that are complex conjugates. In a more general scenario, if the form have complex eigenvalues, the tensor can be diagonalized in terms of complex pairs, where each pair consists of eigenvalues that are complex conjugates of each other. However, to maintain the requirement of having one timelike and three spacelike eigenvalues, our tensor can only admit one pair of complex conjugate eigenvalues [Landau and Lifschitz, 1975].
${K{{}^{3}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&-\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (34)
@@
It should be noted that the only difference between $K^{2}$ and $K^{3}$ could be described via a factor $q$. We choose to deal simultaneously with forms $K^{2}$ and $K^{3}$ with the usage of the parameter $q=\pm 1$ that gives us the 2nd and 3rd forms for $+1$ and $-1$ accordingly.

### v2

@@ 4.2 Canonical forms
Obtaining the canonical forms of a symmetric 2nd-rank tensor can be a challenging task when approached algebraically. In general, the canonical forms contain the minimum number of independent scalars which compose the eigenvalues during the diagonalization [1]. The equation for diagonalization in any 2-rank tensor takes the following form.
${K^{\mu}}_{\nu}x^{\nu}-\lambda x^{\mu}=0$ (29)
In this manner, it becomes clear that the operation of our tensor or the components of our linear vector function on a vector leaves it unchanged. In this context, it is known that every 2-rank symmetric tensor defines a linear mapping that transforms a vector x into another vector v unless the vector is an eigenvector. Equivalently, when a linear vector function satifies $K(\textbf{x})=\lambda\textbf{x}$ ($\mu$ is scalar) then the direction of k is called invariant under the action of the function $K$. We should denote that we are searching for these kinds of directions, which, of course, do not alter the quadratic form of the metric.
@@
In our formalism, the norm of a vector must remain invariant under the action of the linear vector function $K$. Algebraically, the result is that a symmetric matrix $K$ is reducible under an orthogonal transformation with a matrix P to a canonical form $PKP^{-1}$ in which all non-diagonal elements are zero.
@@
A similar analysis can also be found in Landau-Lifshitz’s book on p. 271 regarding the stress-energy-momentum tensor [32]. In this section, the authors mention that this procedure is, in fact, the application of the “Principal Axis Theorem” or “Spectral Theorem” for a matrix. However, it appears that the authors apply diagonalization to a symmetric tensor in a covariant form. This approach is not correct in general. Diagonalization should be applied to a tensor in a mixed tensor form, as shown in equation (29). The latter is always valid because “a term proportional to the metric merely shifts all eigenvalues by the same amount” as denoted in [23].
4.2.1 The presence of null vectors within invariable planes
In line with the work of Churchill, we obtained the canonical forms of Killing tensor [1]. We advise the reader to consult not only Churchill’s and Rainich’s work [2] but also the work of Landau and Lifshitz (p. 271) [32] for a better understanding of the procedure.
In this section, we provide the corresponding relations regarding our formalism, but we do not present the exact same steps as were conducted in [1]. Churchill’s study serves as a sequel to that of Rainich’s concerning a geometrical approach to find the orthogonal directions that determined by the linear symmetric vector function. Both of these works were conducted in a pseudo-Euclidean spacetime in the framework of standard formalism with signature (,+,+,+)(-,+,+,+)( - , + , + , + ).
In this section we focus on the corresponding components of the symmetric linear vector function $K$ which were defined by the way the function transforms the coordinate vectors and they equivalently corresponds to the components of the abstract mixed Killing tensor of second rank $K{{}_{\mu}}^{\nu}$ in the framework of Newman-Penrose formalism with signature (+,,,)(+,-,-,-)( + , - , - , - ).
$K(\boldsymbol{i})=K{{}_{1}}^{1}\boldsymbol{i}+K{{}_{1}}^{2}\boldsymbol{j}+K{{}_{1}}^{3}\boldsymbol{k}+K{{}_{1}}^{4}\boldsymbol{l}$
$K(\boldsymbol{j})=K{{}_{2}}^{1}\boldsymbol{i}+K{{}_{2}}^{2}\boldsymbol{j}+K{{}_{2}}^{3}\boldsymbol{k}+K{{}_{2}}^{4}\boldsymbol{l}$ (31)
$K(\boldsymbol{k})=K{{}_{3}}^{1}\boldsymbol{i}+K{{}_{3}}^{2}\boldsymbol{j}+K{{}_{3}}^{3}\boldsymbol{k}+K{{}_{3}}^{4}\boldsymbol{l}$
$K(\boldsymbol{l})=K{{}_{4}}^{1}\boldsymbol{i}+K{{}_{4}}^{2}\boldsymbol{j}+K{{}_{4}}^{3}\boldsymbol{k}+K{{}_{4}}^{4}\boldsymbol{l}$
It should be noted that due to the symmetry properties (for example $K(\boldsymbol{i})\boldsymbol{j}=K(\boldsymbol{j})\boldsymbol{i}$, etc. ) there are simplifications to the components of our linear symmetric vector function $K$ such as $K{{}_{1}}^{2}=-K{{}_{2}}^{1}$. Next, we present Churchill’s unitary basis vectors with respect to the null tetrads. The first basis vector behaves as a timelike vector (with a negative squared norm), while the rest behave as spacelike vectors (with positive squared norms). The latter can be verified using the orthogonality properties (see relations 4).
$\boldsymbol{i}\equiv\frac{l_{\mu}-n_{\mu}}{\sqrt{2}}dx^{\mu}$
$\boldsymbol{j}\equiv\frac{l_{\mu}+n_{\mu}}{\sqrt{2}}dx^{\mu}$ (32)
$\boldsymbol{k}\equiv i\frac{m_{\mu}-\bar{m}_{\mu}}{\sqrt{2}}dx^{\mu}$
$\boldsymbol{l}\equiv\frac{m_{\mu}+\bar{m}_{\mu}}{\sqrt{2}}dx^{\mu}$
Churchill initiated his study based on the theorem of Rainich who studied the anti-symmetricity of the electromagnetic tensor and he postulated that: “It is known that every linear vector function in four-dimensional space has at least one invariable plane” [2]. Afterwards, in a Lorentzian spacetime there are 3 kinds of 2 dimensional planes. The first case is considered as a singular case and the last two cases are considered as non-singular cases which are also be studied by Landau and Lifshitz in p. 271 [32].
  1. 1.
    A plane with a single null vector, like $\boldsymbol{i}+\boldsymbol{j},\boldsymbol{k}$ plane (Singular case).
  2. 2.
    A plane without null vectors, like $\boldsymbol{k},\boldsymbol{l}$ plane (Non-singular case).
  3. 3.
    A plane with 2 null vectors, like $\boldsymbol{i},\boldsymbol{j}$ plane (Non-singular case).
4.2.2 $K^{0}$ canonical form
We start our technical discussion with the singular case (Case 0 in our study) following the procedure of page 790 in [1]. As postulated an examination of $\boldsymbol{i}+\boldsymbol{j},\boldsymbol{k}$ plane and $\boldsymbol{i}+\boldsymbol{j},\boldsymbol{l}$ plane reveals that “there is always a spacelike direction with eigenvalue $\omega_{4}$”. In connection to this, Churchill defines this invariable direction as a new vector $\boldsymbol{l}$ where
$K(\boldsymbol{l})=\omega_{4}\boldsymbol{l}$
In the remaining 3 dimensional space which is perpendicular to $\boldsymbol{l}$ direction and satisfies the existence of a single null vector there is a function $\phi(\boldsymbol{x})\equiv K(\boldsymbol{x})-\lambda_{1}\boldsymbol{x}$. The $\phi(\boldsymbol{x})$ has the same invariable directions as $K(\boldsymbol{x})$ but different multipliers. Furthermore it transforms every vector $\boldsymbol{x}$ that belongs to $\boldsymbol{i},\boldsymbol{j},\boldsymbol{k}$ space into a vector of $\boldsymbol{i}+\boldsymbol{j},\boldsymbol{k}$ plane, also it transforms every vector in this plane into a null vector aligned with $\boldsymbol{i}+\boldsymbol{j}$ direction. Finally, after the determination of $\phi(\boldsymbol{x})$ function, Churchill results in function $K(\boldsymbol{x})$ where $\lambda_{1},\omega_{4}$ are the only multipliers of the invariable directions of $\boldsymbol{i}+\boldsymbol{j},\boldsymbol{l}$ accordingly.
We present the canonical form obtained by Churchill (relation (21) in [1]), with a single difference in the second term of the first two relations, which arises due to metric signature. The last statement can also be verified through the symmetry property $K(\boldsymbol{i})\boldsymbol{k}=K(\boldsymbol{k})\boldsymbol{i}$.
$K(\boldsymbol{i})=\lambda_{1}\boldsymbol{i}+\boldsymbol{k}$
$K(\boldsymbol{j})=\lambda_{1}\boldsymbol{j}-\boldsymbol{k}$
$K(\boldsymbol{k})=\boldsymbol{i}+\boldsymbol{j}+\lambda_{1}\boldsymbol{k}$
$K(\boldsymbol{l})=\omega_{4}\boldsymbol{l}$
If we try to express the above relations with respect to null tetrads [^6] For the sake of simplification, we present these relations without indices. we get the following relations
$K(n)=\lambda_{1}n-i((-\bar{m})-(-m))$
$K(l)=\lambda_{1}l$ (33)
$K(-\bar{m})=-il-\lambda_{2}(-\bar{m})-\lambda_{7}(-m)$
$K(-m)=+il-\lambda_{7}(-\bar{m})-\lambda_{2}(-m)$
where we define that $-\lambda_{2}\equiv\frac{\omega_{4}+\lambda_{1}}{2}$ and $-\lambda_{7}\equiv\frac{\omega_{4}-\lambda_{1}}{2}$. In the matrix representation we present the mixed Killing form and with the usage of the metric we can obtain the covariant Killing form apparently.
${K{{}^{0}}}{{}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}&0&\bar{p}&p\\ 0&\lambda_{1}&0&0\\ 0&-p&\lambda_{7}&\lambda_{2}\\ 0&-\bar{p}&\lambda_{2}&\lambda_{7}\end{pmatrix}\Longleftrightarrow{K{{}^{0}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&-p&-\bar{p}\\ \lambda_{1}&0&0&0\\ -p&0&\lambda_{7}&\lambda_{2}\\ -\bar{p}&0&\lambda_{2}&\lambda_{7}\end{pmatrix};\hskip 8.5359ptp=-\bar{p}=\pm i$ (34)
Diagonalizing the $K^{0}{{}_{\mu}}^{\nu}$ form one can easily find that this form has one triple eigenvalue, namely $\lambda_{1}=-(\lambda_{2}-\lambda_{7})$ which is placed in a 3-dimensional Jordan canonical form. In addition, the other eigenvalue described by the form $\omega_{4}=-(\lambda_{2}+\lambda_{7})$.
4.2.3 $K^{1}$, $K^{2}$, $K^{3}$ canonical forms
In this section we are going to provide the general form of the mixed Killing tensor concerning all three cases. Afterwards, we will operate a classification based on the sign of the discriminant of the equation (36) that emerges during the diagonalization.
Starting by the assumption of existence of zero null vectors in the invariable plane of $K(\boldsymbol{x})$ we can find new $\boldsymbol{k}$ and $\boldsymbol{l}$ vectors lie in this invariable plane and transform them to invariable directions. Building upon this, the new $\boldsymbol{i},\boldsymbol{j}$ plane is also invariable and perpendicular to the former.
On the other hand initiating by the assumption of existence of two null vectors in the invariable plane of $K(\boldsymbol{x})$ we can find new $\boldsymbol{i},\boldsymbol{j}$ which can be made to fall into this invariable plane. Using the symmetry property we can find new $\boldsymbol{k}$ and $\boldsymbol{l}$ vectors which are lie to an invariable plane with zero null vectors, therefore in this plane there are two invariable directions as in the previous non-singular case.
Hence, both cases yield the same 2-rank mixed Killing tensor, with the main difference being the number of null vectors in the $\boldsymbol{i},\boldsymbol{j}$ plane.
$K(\boldsymbol{i})=\alpha\boldsymbol{i}+\beta\boldsymbol{j}$
$K(\boldsymbol{j})=-\beta\boldsymbol{i}+\gamma\boldsymbol{j}$ (35)
$K(\boldsymbol{k})=\omega_{3}\boldsymbol{k}$
$K(\boldsymbol{l})=\omega_{4}\boldsymbol{l}$
An examination of this plane will distinct the different cases of the canonical forms. A vector $\boldsymbol{x}=x^{1}\boldsymbol{i}+x^{2}\boldsymbol{j}$ belongs to invariable direction if
$K(x^{1}\boldsymbol{i}+x^{2}\boldsymbol{j})=\omega(x^{1}\boldsymbol{i}+x^{2}\boldsymbol{j})$
Subsequently, we take
$x^{1}(\alpha-\omega)-x^{2}\beta=0$
$x^{1}\beta+x^{2}(\gamma-\omega)=0$
By eliminating $x^{1}$ and $x^{2}$, we investigate solutions where neither of these variables are zero.
$\omega^{2}-\omega(\alpha+\gamma)+\beta^{2}+\alpha\gamma=0\Longrightarrow\omega_{1,2}=\frac{(\alpha+\gamma)\pm\sqrt{(\alpha+\gamma)^{2}-4(\beta^{2}+\alpha\gamma)}}{2}$ (36)
This equation yield two solutions $\omega_{1}$ and $\omega_{2}$. Within this scheme, three cases arise from the examination of the discriminant. The first case (Case 1) emerges when the discriminant is set to zero, resulting in a double root with $\omega_{1}=\omega_{2}=\frac{\alpha+\gamma}{2}$. In contrast to Churchill’s study during this examination we did not consider the cases where 1) $\beta=0$ as it is a special case of Case 3 and 2) $\beta\neq 0$, where $\omega_{1}=\omega_{2}=0$. However, in Case 1 we obtained two different forms that result in the same diagonalized form.
This is evident if we use the relation (14) from [1], where $\psi_{1}=2\sqrt{\psi_{2}}=\omega_{1}$.
${K{{}^{1a}}}{{}_{\mu}}^{\nu}=\begin{pmatrix}0&\lambda_{1}&0&0\\ -\lambda_{1}&\lambda_{0}&0&0\\ 0&0&\omega_{3}&0\\ 0&0&0&\omega_{4}\end{pmatrix}\hskip 2.84544pt,{K{{}^{1b}}}{{}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ -\lambda_{1}&0&0&0\\ 0&0&\omega_{3}&0\\ 0&0&0&\omega_{4}\end{pmatrix}$ (37)
The transformation in null basis vectors using the relations (32) and setting $\omega_{1}=\lambda_{1}$, $\lambda_{0}=\pm 2\omega_{1}$, $-\lambda_{2}\equiv\omega_{4}+\omega_{3}$ and $-\lambda_{7}\equiv\omega_{4}-\omega_{3}$  [^7] It should be noted that $\lambda_{0}$ does not contribute to the calculations and one could verify that for any scalar value of $\lambda_{0}$ the diagonalization results in a 2 dimensional Jordan block with a double eigenvalue $\lambda_{1}$ in the upper-left block. yields the two canonical forms with one double eigenvalue $\lambda_{1}$ and two single eigenvalues $-(\lambda_{2}\pm\lambda_{7})$.
${K{{}^{1a}}}_{\mu\nu}=\begin{pmatrix}0&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}\hskip 2.84544pt,{K{{}^{1b}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&0&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (38)
@@ 4.2.3 $K^{1}$, $K^{2}$, $K^{3}$ canonical forms
This case produces two forms but considered as one by us. The reason is based on the concept of the symmetrical null tetrad frame which introduced by Debever in [33]. This symmetry refers to the interchanges between the tetrads $n^{\mu}\leftrightarrow l^{\mu}$ and $m^{\mu}\leftrightarrow\bar{m}^{\mu}$.
The next case is the most general one, as it contains two distinct real eigenvalues and corresponds to Case 2. This case arises when the discriminant is positive. Hence, there are 4 invariable directions, and the mixed Killing tensor appears in a diagonal form with eigenvalues $\omega_{1},\omega_{2},\omega_{3},\omega_{4}$. Using again the relations (32) and setting $\lambda_{1}\equiv\omega_{2}+\omega_{1}$ and $\lambda_{0}\equiv\omega_{2}-\omega_{1}$, $-\lambda_{2}\equiv\omega_{4}+\omega_{3}$ and $-\lambda_{7}\equiv\omega_{4}-\omega_{3}$
${K{{}^{2}}}{{}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}&\lambda_{0}&0&0\\ \lambda_{0}&\lambda_{1}&0&0\\ 0&0&-\lambda_{2}&-\lambda_{7}\\ 0&0&-\lambda_{7}&-\lambda_{2}\end{pmatrix}\hskip 2.84544pt\Longleftrightarrow{K{{}^{2}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (39)
Finally, Case 3 arises when the discriminant takes negative values. In this scenario, there are no invariable directions within the $\boldsymbol{i},\boldsymbol{j}$ plane. Consequently, we are unable to find more useful directions and must work with the available one. Taking advantage of relations (35) and combine them with relations (32) we get the following
$K(n)=\frac{\alpha+\gamma}{2}n+\frac{\alpha-\gamma-2\beta}{2}l$ (40)
$K(l)=-\frac{\alpha-\gamma-2\beta}{2}n+\frac{\alpha+\gamma}{2}l$ (41)
In this point we define $\lambda_{1}\equiv\frac{\alpha+\gamma}{2}$, $\lambda_{0}\equiv\frac{\alpha-\gamma-2\beta}{2}$, $-\lambda_{2}\equiv\omega_{4}+\omega_{3}$ and $-\lambda_{7}\equiv\omega_{4}-\omega_{3}$
It should be noted that in this manner the existence of one timelike and one spacelike eigenvector is satisfied since it serves as a constraint to us. Generally speaking this is the reason why in a pseduo-Euclidean spacetime we cannot have two pairs of complex conjugate eigenvalues [32]. After the definitions above we result in the mixed tensor form and with the usage of metric of our formalism we obtain the covariant one.
${K{{}^{3}}}{{}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}&\lambda_{0}&0&0\\ -\lambda_{0}&\lambda_{1}&0&0\\ 0&0&-\lambda_{2}&-\lambda_{7}\\ 0&0&-\lambda_{7}&-\lambda_{2}\end{pmatrix}\hskip 2.84544pt\Longleftrightarrow{K{{}^{3}}}_{\mu\nu}=\begin{pmatrix}\lambda_{0}&\lambda_{1}&0&0\\ \lambda_{1}&-\lambda_{0}&0&0\\ 0&0&\lambda_{7}&\lambda_{2}\\ 0&0&\lambda_{2}&\lambda_{7}\end{pmatrix}$ (42)
@@
It should be noted that the only difference between $K^{1}$, $K^{2}$ and $K^{3}$ could be described via a factor $q\equiv 0,\pm 1$.